# Cryptographic Characterization of Quantum Advantage

### v1

@@ Abstract
Quantum computational advantage refers to an existence of computational tasks that are easy for quantum computing but hard for classical one. Unconditionally showing quantum advantage is beyond our current understanding of complexity theory, and therefore some computational assumptions are needed. Which complexity assumption is necessary and sufficient for quantum advantage? In this paper, we show that inefficient-verifier proofs of quantumness (IV-PoQ) exist if and only if classically-secure one-way puzzles (OWPuzzs) exist. As far as we know, this is the first time that a complete cryptographic characterization of quantum advantage is obtained. IV-PoQ [Morimae and Yamakawa 2024] are a variant of proofs of quantumness (PoQ) [Brakerski et al. 2021] where the verifier is efficient during the interaction but may use unbounded time afterward. IV-PoQ capture various types of quantum advantage previously studied, such as sampling advantage and searching advantage. Previous work [Morimae and Yamakawa 2024] showed that IV-PoQ can be constructed from OWFs, but a construction of IV-PoQ from weaker assumptions was left open. Our result solves the open problem, because OWPuzzs are believed to be weaker than OWFs. OWPuzzs [Khurana and Tomer 2024] are one of the most fundamental quantum cryptographic primitives implied by many quantum cryptographic primitives weaker than one-way functions (OWFs), such as pseudorandom unitaries (PRUs), pseudorandom state generators (PRSGs), and one-way state generators (OWSGs). The equivalence between IV-PoQ and classically-secure OWPuzzs therefore highlights that if there is no quantum advantage, then these fundamental primitives do not exist. The equivalence also means that quantum advantage is an example of the applications of OWPuzzs. Except for commitments, no application of OWPuzzs was known before. Our result shows that quantum advantage is another application of OWPuzzs. Moreover, it is the first quantum computation classical communication (QCCC) application of OWPuzzs. To show the main result, we introduce several new concepts and show some results that will be of independent interest. In particular, we introduce an interactive (and average-case) version of sampling problems where the task is to sample the transcript obtained by a classical interaction between two quantum polynomial-time algorithms. We show that quantum advantage in interactive sampling problems is equivalent to the existence of IV-PoQ, which is considered as an interactive (and average-case) version of Aaronson’s result [Aaronson 2014], $\mathbf{SampBQP}\neq\mathbf{SampBPP}\Leftrightarrow\mathbf{FBQP}\neq\mathbf{FBPP}$. Finally, we also introduce zero-knowledge IV-PoQ and study sufficient and necessary conditions for their existence.
@@ 1 Introduction
Quantum computational advantage refers to the existence of computational tasks that are easy for quantum computing but hard for classical one. Unconditionally showing quantum advantage is extremely hard, and is beyond our current understanding of complexity theory. [^1] There are several interesting results (such as [BGK18]) that show unconditional quantum advantage by restricting classical computing. In this paper, we consider any polynomial-time classical computing. Some computational assumptions are therefore required. Which complexity assumption is necessary and sufficient for quantum advantage? As far as we know, no complete characterization of quantum advantage was obtained before.
@@
IV-PoQ [MY24] are the same as PoQ except that the verifier’s final computation to make the decision can be unbounded. IV-PoQ are a generalization of PoQ, and as we will explain later, IV-PoQ capture various types of quantum advantage studied so far, such as sampling advantage and searching advantage.
@@
In classical cryptography, the existence of OWFs is the minimum assumption [IL89], because many primitives exist if and only if OWFs exist, such as pseudorandom generators (PRGs), pseudorandom functions (PRFs), zero-knowledge, commitments, digital signatures, and secret-key encryptions (SKE), and almost all primitives imply OWFs. On the other hand, recent active studies have revealed that in quantum cryptography OWFs are not necessarily the minimum assumption. Many fundamental primitives have been introduced, such as pseudorandom unitaries (PRUs) [JLS18], pseudorandom function-like state generators (PRFSGs) [AQY22], unpredictable state generators (UPSGs) [MYY24], pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22], EFI pairs [BCQ23], and one-way puzzles (OWPuzzs) [KT24a]. They could exist even if OWFs do not exist [Kre21, KQST23, LMW24], but still imply several useful applications such as message authentication codes [AQY22], commitments [MY22, AQY22], multi-party computations [MY22, AQY22], secret-key encryptions [AQY22], private-key quantum money [JLS18], digital signatures [MY22], etc.
@@
  • several quantum computation classical communication (QCCC) primitives [CGG24a, KT24a],
@@
On the other hand, although many primitives imply OWPuzzs, no application of OWPuzzs is known except for commitments [KT24a]. Finding more applications of OWPuzzs is one of the most important goals in this field. Our result Theorem 1.1 shows that OWPuzzs imply quantum advantage, which demonstrates that quantum advantage is another application of OWPuzzs. Moreover, we emphasize that this is the first application of OWPuzzs in the QCCC setting: IV-PoQ are a QCCC primitive because the communication between the verifier and the prover is classical, while commitments [KT24a] constructed from OWPuzzs are those over quantum channels.
@@
In addition to (IV-)PoQ, there are mainly two other approaches to demonstrate quantum advantage, namely searching and sampling. Here we argue that IV-PoQ capture both of them, and therefore identifying a necessary and sufficient assumption for the existence of IV-PoQ is significant.
A sampling problem is a task of sampling from some distributions. There are several distributions that are easy to sample with QPT algorithms but hard with PPT algorithms, such as output ditributions of random quantum circuits [BFNV19], Boson Sampling circuits [AA11], constant-depth circuits [TD04], IQP circuits [BJS11, BMS16], and one-clean-qubit circuits [FKM+18]. Several assumptions are known to be sufficient for quantum advantage in sampling problems, but these assumptions are newly-introduced assumptions that were not studied before such as an average-case #$\#{\bf P}$-hardness of approximating some functions. Moreover, quantum advantage in sampling problems is in general not known to be verifiable (even inefficiently). On the other hand, one advantage of sampling problems (and other problems relying on newly-introduced assumptions) is that experimental realizations with NISQ machines seem to be easier. [^7] Although NISQ experimental realizations of quantum advantage are very important goals, in this paper, we focus on theoretical upper and lower bounds by assuming that any polynomial-time quantum computing is possible. As we will explain later, we introduce an average-case version of $\mathbf{SampBQP}\neq\mathbf{SampBPP}$, and show that such an average-case version of sampling advantage is equivalent to the existence of non-interactive IV-PoQ. IV-PoQ therefore capture sampling advantage.
A search problem is a task of finding an element $z$ that satisfies a relation $R(z)=1$. Several search problems have been shown to be easy for QPT algorithms but hard for PPT algorithms. Their classical hardness, however, relies on newly-introduced assumptions that were not studied before, such as QUATH [AC17] and XQUATH [AG19], or relies on random oracles [Aar10, ACC+23]. One advantage of these search problems over sampling problems is that quantum advantage can be verified at least inefficiently when $R$ is computable. (We can check $R(z)=1$ or not by computing $R(z)$.) Because such inefficiently-verifiable search advantage is equivalent to the existence of non-interactive IV-PoQ, IV-PoQ capture inefficiently-verifiable search problems. There are some search problems that are efficiently verifiable such as Factoring [Sho94] and Yamakawa-Zhandry problem [YZ24] but the former is based on the hardness of a specific problem, and the latter relies on the random oracle model. Quantum advantage in efficiently-verifiable search problems is captured by non-interactive PoQ, and therefore by IV-PoQ.
@@
In summary, we have shown that IV-PoQ exist if and only if classically-secure OWPuzzs exist. We believe that this result is significant mainly because of the following five reasons.
@@
    As far as we know, this is the first time to give a complete cryptographic characterization of quantum advantage.
@@
    IV-PoQ capture various types of quantum advantage studied so far including sampling advantage and searching advantage.
@@
    No application of OWPuzzs was known before except for commitments (and therefore multiparty computations). We show that quantum advantage is another application of OWPuzzs. Moreover, it is the first QCCC application of OWPuzzs.
QAS/OWFcs-OWPuzzsqs-OWPuzzsOWSGsEFIDEFIPRSGs1-PRSGsPRFSGsPRUsqs-OWFscs-OWFsIV-PoQpublic-coin IV-PoQquantum-verifier IV-PoQnon-interactive IV-PoQhonest-verifier statistical zero-knowledge IV-PoQcomputational zero-knowledge IV-PoQInt-QASsQASsqs-OWFs or $\mathbf{SampBPP}\neq\mathbf{SampBQP}$
Figure 1: Summary of results. Black arrows are known or trivial implications. Red arrows are our results. “qs” stands for quantumly-secure and “cs” stands for classically-secure. 1-PRSGs are single-copy-secure PRSGs [MY22].
@@ 1.1 Additional Results
We show the following result. [^8] Note that “quantumly-secure” in the theorem is not a typo. The reason why we can get quantumly-secure OWFs from classically-secure OWPuzzs is, roughly speaking, that if $\mathbf{SampBPP}=\mathbf{SampBQP}$, then classically-secure OWFs means quantumly-secure OWFs. For details, see Section 4.
@@
Intuitively, Equation 1 means that the output distribution of the QPT algorithm $\mathcal{A}$ cannot be classically efficiently sampled. How is this different from the more studied notion, $\mathbf{SampBPP}\neq\mathbf{SampBQP}$? The existence of QASs can be considered as an average-case version of $\mathbf{SampBPP}\neq\mathbf{SampBQP}$101010 $\mathbf{SampBPP}$ and $\mathbf{SampBQP}$ are worst-case complexity classes, and therefore $\{\mathcal{A}(1^{\lambda})\}_{\lambda}\not\in\mathbf{SampBPP}$ means that there is at least one $\lambda\in\mathbb{N}$ such that $\mathcal{A}(1^{\lambda})$ cannot be classically efficiently sampled, while the definition of QASs requires that $\{\mathcal{A}(1^{\lambda})\}_{\lambda}$ cannot be classically efficiently sampled for all sufficiently large $\lambda\in\mathbb{N}$. In addition, there is a subtle technical difference: $\{\mathcal{A}_{\lambda}\}_{\lambda}\not\in\mathbf{SampBPP}$ means that $\mathcal{A}(1^{\lambda})$ cannot be classically efficiently sampled for a precision $\epsilon$, but this $\epsilon$ could be ${\mathsf{negl}}(\lambda)$, while QASs requires that the precision is $1/{\mathrm{poly}}(\lambda)$. For more details, see Section 3. In fact, the existence of QASs implies $\mathbf{SampBPP}\neq\mathbf{SampBQP}$ (Lemma 3.4), but the inverse does not seem to hold. In order to show our main result, the worst-case notion of $\mathbf{SampBPP}\neq\mathbf{SampBQP}$ is not enough for our cryptographic (and therefore average-case) argument, and therefore we introduce QASs. We believe that the new concept, QASs, will be useful for other future studies of quantum sampling advantage in the context of quantum cryptography.
@@ 2.2 One-Way Functions and Distributional One-Way Functions
It is well-known that OWFs exist if and only if DistOWFs exist [IL89, Lemma 1].161616[IL89] only provides a proof sketch. Its full proof can be found in [Imp92, Theorem 4.2.2] By inspecting its proof, one can see that the proof gives a “security-parameter-wise” reduction, i.e., if the base DistOWF is secure on inputs of length $\lambda$, then the resulting OWF is secure on inputs of length $n(\lambda)$ for some efficiently computable polynomial $n$ [^17] More precisely, for any efficiently computable function $f:\{0,1\}^{*}\rightarrow\{0,1\}^{*}$ and polynomial $p$, there are an efficiently computable function $g:\{0,1\}^{*}\rightarrow\{0,1\}^{*}$ and efficiently computable polynomial $n$ that satisfy the following: For any PPT adversary $\mathcal{A}$ and a polynomial $q$, there is a PPT adversary $\mathcal{B}$ such that for any $\lambda\in\mathbb{N}$, if $\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{n(\lambda)},x^{\prime}\leftarrow\mathcal{A}(1^{n(\lambda)},g(x))]>\frac{1}{q(\lambda)},$ (7) then $\mathsf{SD}(\{(x,f(x))\}_{x\leftarrow\{0,1\}^{\lambda}},\{(\mathcal{B}(1^{\lambda},f(x)),f(x))\}_{x\leftarrow\{0,1\}^{\lambda}})<\frac{1}{p(\lambda)}.$ (8) This implies equivalence between OWFs on $\Sigma$ and DistOWFs on $\Sigma$ on any subset $\Sigma\subseteq\mathbb{N}$.

### v2

@@ Abstract
Quantum computational advantage refers to an existence of computational tasks that are easy for quantum computing but hard for classical one. Unconditionally showing quantum advantage is beyond our current understanding of complexity theory, and therefore some computational assumptions are needed. Which complexity assumption is necessary and sufficient for quantum advantage? In this paper, we show that inefficient-verifier proofs of quantumness (IV-PoQ) exist if and only if classically-secure one-way puzzles (OWPuzzs) exist. As far as we know, this is the first time that a complete cryptographic characterization of quantum advantage is obtained. IV-PoQ are a generalization of proofs of quantumness (PoQ) where the verifier is efficient during the interaction but may use unbounded time afterward. IV-PoQ capture various types of quantum advantage previously studied, such as sampling-based quantum advantage and searching-based one. Previous work [Morimae and Yamakawa, Crypto 2024] showed that IV-PoQ can be constructed from OWFs, but a construction of IV-PoQ from weaker assumptions was left open. Our result solves the open problem, because OWPuzzs are believed to be weaker than OWFs. OWPuzzs are one of the most fundamental quantum cryptographic primitives implied by many quantum cryptographic primitives weaker than one-way functions (OWFs), such as pseudorandom unitaries (PRUs), pseudorandom state generators (PRSGs), and one-way state generators (OWSGs). The equivalence between IV-PoQ and classically-secure OWPuzzs therefore highlights that if there is no quantum advantage, then these fundamental cryptographic primitives do not exist. The equivalence also means that quantum advantage is an example of the applications of OWPuzzs. Except for commitments, no application of OWPuzzs was known before. Our result shows that quantum advantage is another application of OWPuzzs, which solves the open question of [Chung, Goldin, and Gray, Crypto 2024]. Moreover, it is the first quantum-computation-classical-communication (QCCC) application of OWPuzzs. To show the main result, we introduce several new concepts and show some results that will be of independent interest. In particular, we introduce an interactive (and average-case) version of sampling problems where the task is to sample the transcript obtained by a classical interaction between two quantum polynomial-time algorithms. We show that quantum advantage in interactive sampling problems is equivalent to the existence of IV-PoQ, which is considered as an interactive (and average-case) version of Aaronson’s result [Aaronson, TCS 2014], $\mathbf{SampBQP}\neq\mathbf{SampBPP}\Leftrightarrow\mathbf{FBQP}\neq\mathbf{FBPP}$. Finally, we also introduce zero-knowledge IV-PoQ and study sufficient and necessary conditions for their existence.
@@ 1 Introduction
Quantum computational advantage refers to the existence of computational tasks that are easy for quantum computing but hard for classical one. Unconditionally showing quantum advantage is extremely hard, and is beyond our current understanding of complexity theory. [^1] There are several interesting results (such as [BGK18]) that show unconditional quantum advantage by restricting classical computing. In this paper, we consider any polynomial-time classical computing. Some computational assumptions are therefore required. Which complexity assumption is necessary and sufficient for quantum advantage? As far as we know, no complete characterization of quantum advantage has been achieved before.
@@
IV-PoQ [MY24] are the same as PoQ except that the verifier’s final computation to make the decision can be unbounded. IV-PoQ are a generalization of PoQ, and as we will explain later, IV-PoQ capture various types of quantum advantage studied so far, such as sampling-based quantum advantage and searching-based one.
@@
In classical cryptography, the existence of OWFs is the minimum assumption [IL89], because many primitives exist if and only if OWFs exist, such as pseudorandom generators (PRGs), pseudorandom functions (PRFs), zero-knowledge, commitments, digital signatures, and secret-key encryptions (SKE), and almost all primitives imply OWFs. On the other hand, recent active studies have revealed that in quantum cryptography OWFs are not necessarily the minimum assumption. Many fundamental primitives have been introduced, such as pseudorandom unitaries (PRUs) [JLS18], pseudorandom function-like state generators (PRFSGs) [AQY22], unpredictable state generators (UPSGs) [MYY24], pseudorandom state generators (PRSGs) [JLS18], one-way state generators (OWSGs) [MY22], EFI pairs [BCQ23], and one-way puzzles (OWPuzzs) [KT24a]. They could exist even if OWFs do not exist [Kre21, KQST23, LMW24], but still imply several useful applications such as message authentication codes [AQY22], commitments [MY22, AQY22], multi-party computations [MY22, AQY22, BCKM21, GLSV21], secret-key encryptions [AQY22], private-key quantum money [JLS18], digital signatures [MY22], etc.
@@
  • many quantum-computation-classical-communication (QCCC) primitives [^6] Here, QCCC primitives are primitives with local quantum computation and classical communication. For example, in QCCC commitments, sender and receiver are QPT, while the message exchanged between them are classical. [CGG24, KT24a, GMMY24],
@@
On the other hand, although many primitives imply OWPuzzs, no application of OWPuzzs is known except for commitments [KT24a]. Finding more applications of OWPuzzs is one of the most important goals in this field. Our result Theorem 1.1 shows that OWPuzzs imply quantum advantage, which demonstrates that quantum advantage is another application of OWPuzzs. Moreover, we emphasize that this is the first application of OWPuzzs in the QCCC setting: IV-PoQ are a QCCC primitive because the communication between the verifier and the prover is classical, while commitments [KT24a] constructed from OWPuzzs are those over quantum channels. The question of the existence of QCCC applications of OWPuzzs was raised in [CGG24]. We solve the open problem.
@@
In addition to (IV-)PoQ, there are mainly two other approaches to demonstrate quantum advantage, namely sampling-based quantum advantage and searching-based one. Here we argue that IV-PoQ capture both of them, and therefore identifying a necessary and sufficient assumption for the existence of IV-PoQ is significant.
A sampling problem is a task of sampling from some distributions. There are several distributions that are easy to sample with QPT algorithms but hard with PPT algorithms, such as output distributions of random quantum circuits [BFNV19], Boson Sampling circuits [AA11], constant-depth circuits [TD04], IQP circuits [BJS11, BMS16], and one-clean-qubit circuits [FKM+18, Mor17]. Several assumptions are known to be sufficient for quantum advantage in sampling problems, but these assumptions are newly-introduced assumptions that were not studied before such as an average-case #$\#{\bf P}$-hardness of approximating some functions. Moreover, quantum advantage in sampling problems is in general not known to be verifiable (even inefficiently). On the other hand, one advantage of sampling-based quantum advantage (and others relying on newly-introduced assumptions) is that experimental realizations with NISQ machines seem to be easier. [^8] Although NISQ experimental realizations of quantum advantage are very important goals, in this paper, we focus on theoretical upper and lower bounds by assuming that any polynomial-time quantum computing is possible. As we will explain later, we introduce an average-case version of $\mathbf{SampBQP}\neq\mathbf{SampBPP}$ [^9] For the definitions of $\mathbf{SampBQP}$ and $\mathbf{SampBPP}$, see Definitions 2.23 and 2.24., and show that it is equivalent to the existence of non-interactive IV-PoQ. IV-PoQ therefore capture sampling-based quantum advantage.
A search problem is a task of finding an element $z$ that satisfies a relation $R(z)=1$. Several search problems have been shown to be easy for QPT algorithms but hard for PPT algorithms. Their classical hardness, however, relies on newly-introduced assumptions that were not studied before, such as QUATH [AC17] and XQUATH [AG19], or relies on random oracles [Aar10, ACC+23]. One advantage of searching-based quantum advantage over sampling-based one is that quantum advantage can be verified at least inefficiently when $R$ is computable. (We can check $R(z)=1$ or not by computing $R(z)$.) Because such inefficiently-verifiable searching-based quantum advantage is equivalent to the existence of non-interactive IV-PoQ, IV-PoQ capture inefficiently-verifiable searching-based quantum advantage. There are some search problems that are efficiently verifiable such as Factoring [Sho94] and Yamakawa-Zhandry problem [YZ24] but the former is based on the hardness of a specific problem, and the latter relies on the random oracle model. Quantum advantage in efficiently-verifiable search problems is captured by non-interactive PoQ, and therefore by IV-PoQ.
@@
In summary, we have shown that IV-PoQ are existentially equivalent to classically-secure OWPuzzs. We believe that this result is significant mainly because of the following five reasons.
@@
    As far as we know, this is the first time that a complete cryptographic characterization of quantum advantage is achieved.
@@
    IV-PoQ capture various types of quantum advantage studied so far including sampling-based quantum advantage and searching-based one.
@@
    No application of OWPuzzs was known before except for commitments (and therefore multiparty computations). We show that quantum advantage is another application of OWPuzzs. Moreover, it is the first QCCC application of OWPuzzs. This solves the open problem of [CGG24].
QAS/OWFcs-OWPuzzsqs-OWPuzzsOWSGsEFIDEFIPRSGs1-PRSGsPRFSGsPRUsqs-OWFscs-OWFsIV-PoQpublic-coin IV-PoQquantum-verifier IV-PoQnon-interactive IV-PoQhonest-verifier statistical zero-knowledge IV-PoQcomputational zero-knowledge IV-PoQInt-QASsQASsqs-OWFs or $\mathbf{SampBPP}\neq\mathbf{SampBQP}$$\mathbf{SampBPP}\neq\mathbf{SampBQP}$QAA + $\mathbf{P}^{\mathbf{\#P}}\not\subseteq\mathbf{(io)BPP}^{\mathbf{NP}}$
Figure 1: Summary of results. Black arrows are known or trivial implications. Red arrows are our results. “qs” stands for quantumly-secure and “cs” stands for classically-secure. 1-PRSGs are single-copy-secure PRSGs [MY22]. QAA stands for the quantum advantage assumption.
@@ 1.1 Additional Results
In the field of quantum advantage, more studied notion of sampling-based quantum advantage is $\mathbf{SampBQP}\neq\mathbf{SampBPP}$. We can show the following relation between IV-PoQ and $\mathbf{SampBQP}\neq\mathbf{SampBPP}$: [^10] Note that “quantumly-secure” in the theorem is not a typo. The reason why we can get quantumly-secure OWFs from classically-secure OWPuzzs is, roughly speaking, that if $\mathbf{SampBPP}=\mathbf{SampBQP}$, then classically-secure OWFs means quantumly-secure OWFs. For details, see Section 4.
@@
Intuitively, Equation 1 means that the output distribution of the QPT algorithm $\mathcal{A}$ cannot be classically efficiently sampled. How is this different from the more studied notion, $\mathbf{SampBPP}\neq\mathbf{SampBQP}$? The existence of QASs can be considered as an average-case version of $\mathbf{SampBPP}\neq\mathbf{SampBQP}$121212 $\mathbf{SampBPP}$ and $\mathbf{SampBQP}$ are worst-case complexity classes, and therefore $\{\mathcal{A}(1^{\lambda})\}_{\lambda}\not\in\mathbf{SampBPP}$ means that there is at least one $\lambda\in\mathbb{N}$ such that $\mathcal{A}(1^{\lambda})$ cannot be classically efficiently sampled, while the definition of QASs requires that $\{\mathcal{A}(1^{\lambda})\}_{\lambda}$ cannot be classically efficiently sampled for all sufficiently large $\lambda\in\mathbb{N}$. In addition, there is a subtle technical difference: $\{\mathcal{A}_{\lambda}\}_{\lambda}\not\in\mathbf{SampBPP}$ means that $\mathcal{A}(1^{\lambda})$ cannot be classically efficiently sampled for a precision $\epsilon$, but this $\epsilon$ could be ${\mathsf{negl}}(\lambda)$, while QASs requires that the precision is $1/{\mathrm{poly}}(\lambda)$. For more details, see Section 3. In fact, the existence of QASs implies $\mathbf{SampBPP}\neq\mathbf{SampBQP}$ (Lemma 3.4), but the inverse does not seem to hold. In order to show our main result, the worst-case notion of $\mathbf{SampBPP}\neq\mathbf{SampBQP}$ is not enough for our cryptographic (and therefore average-case) argument, and therefore we introduce QASs. We believe that the new concept, QASs, will be useful for other future studies of quantum sampling advantage in the context of quantum cryptography.
@@
In previous works on sampling-based quantum advantage [BFNV19, AA11, BMS16, Mor17], $\mathbf{SampBQP}\neq\mathbf{SampBPP}$ was derived based on three assumptions. First one is the complexity assumption of $\mathbf{P}^{\mathbf{\#P}}\not\subseteq\mathbf{BPP}^{\mathbf{NP}}$, which especially means that the polynomial-time hierarchy will not collapse to the third level. Second one is the assumption that computing some functions (such as matrix permanents and Ising partition functions) within a certain multiplicative error is #$\mathbf{\#P}$-hard on average. The third one is so-called the anti-concentration, which roughly means that the output probability of the quantum algorithm is not so concentrating. [^13] For several models, such as the IQP model [BMS16] and the one-clean-qubit model [Mor17], the anti-concentration can be shown. For simplicity, we call the combination of the last two assumptions just quantum advantage assumption.
The quantum advantage assumption has been traditionally studied in the field of quantum advantage to show sampling-based quantum advantage (in terms of $\mathbf{SampBQP}\neq\mathbf{SampBPP}$) of several “non-universal” models such as random quantum circuits [BFNV19], Boson Sampling circuits [AA11], IQP circuits [BMS16], and one-clean-qubit circuits [Mor17]. We can show that the quantum advantage assumption also implies the existence of QASs. [^14] This result was not included in the previous version of this manuscript. We obtained this result after reading [KT24b].
Theorem 1.4.
If the quantum advantage assumption holds and $\mathbf{P}^{\mathbf{\#P}}\not\subseteq\mathbf{ioBPP}^{\mathbf{NP}}$, then QASs exist.
As we have explained above, QASs are cryptographically a more natural notion of sampling-based quantum advantage, and the existence of QASs seems to be stronger than $\mathbf{SampBQP}\neq\mathbf{SampBPP}$. Therefore this theorem reveals that the quantum advantage assumption traditionally studied in the field actually implies a stronger form of quantum advantage (modulo the difference between $\mathbf{BPP}$ and $\mathbf{ioBPP}$).
@@ 2.2 One-Way Functions and Distributional One-Way Functions
It is well-known that OWFs exist if and only if DistOWFs exist [IL89, Lemma 1].202020[IL89] only provides a proof sketch. Its full proof can be found in [Imp92, Theorem 4.2.2] By inspecting its proof, one can see that the proof gives a “security-parameter-wise” reduction, i.e., if the base DistOWF is secure on inputs of length $\lambda$, then the resulting OWF is secure on inputs of length $n(\lambda)$ for some efficiently computable polynomial $n$ [^21] More precisely, for any efficiently computable function $f:\{0,1\}^{*}\rightarrow\{0,1\}^{*}$ and polynomial $p$, there are an efficiently computable function $g:\{0,1\}^{*}\rightarrow\{0,1\}^{*}$ and efficiently computable polynomial $n$ that satisfy the following: For any PPT adversary $\mathcal{A}$ and a polynomial $q$, there is a PPT adversary $\mathcal{B}$ such that for any $\lambda\in\mathbb{N}$, if $\Pr[g(x^{\prime})=g(x):x\leftarrow\{0,1\}^{n(\lambda)},x^{\prime}\leftarrow\mathcal{A}(1^{n(\lambda)},g(x))]>\frac{1}{q(\lambda)},$ (7) then $\mathsf{SD}(\{(x,f(x))\}_{x\leftarrow\{0,1\}^{\lambda}},\{(\mathcal{B}(1^{\lambda},f(x)),f(x))\}_{x\leftarrow\{0,1\}^{\lambda}})<\frac{1}{p(\lambda)}.$ (8) This implies equivalence between OWFs on $\Sigma$ and DistOWFs on $\Sigma$ on any subset $\Sigma\subseteq\mathbb{N}$.